(3.3.7) For a data set of the pulse rates for a sample of adult​ females, the lowest pulse rate is 34 beats per​ minute, the mean of the listed pulse rates is x=74.0 beats per​ minute, and their standard deviation is s=15.8 beats per minute.
a. What is the difference between the pulse rate of 34 beats per minute and the mean pulse rate of the​ females?
b. How many standard deviations is that​ [the difference found in part​ (a)]?
c. Convert the pulse rate of 34 beats per minutes to a z score.
d. If we consider pulse rates that convert to z scores between 2 and 2 to be neither significantly low nor significantly​ high, is the pulse rate of beats per minute​ significant?
1
Expert's answer
2020-11-19T17:38:04-0500

a. The difference between the pulse rate of 34 beats per minute and the mean pulse rate of the​ females= 74-34=40


b. standard deviations is 15.8


c. Z score= "\frac{x-mean}{sd}= \frac{74-34}{15.8}= 2.53"


d. Since 2.53 does lie inside -2 and 2 (range of non-significant values) , thus z score is significant and pulse rate of 34 beats per minute is also significant. =